# Rapid Detection of High-impedance Arc Faults in Medium Voltage Electrical Distribution Systems

### v1

@@ Abstract
High-impedance arc faults in AC power systems have the potential to lead to catastrophic accidents. However, significant challenges exist in identifying these faults because of the much weaker characteristics and variety when grounded with different surfaces. Addressing a noteworthy gap in prior research, which largely focused on arc fault detection in low-voltage systems. A novel approach has been applied that offers rapid arc fault detection for medium voltage distribution lines. In contrast to existing black-box feature-based approaches, Hankel alternative view of the Koopman (HAVOK) analysis developed from nonlinear dynamics has been applied which offers not only interpretable features but also opens up new application options in the area of arc fault detection. The method displays a much faster detection speed within 0.45 ms making it appropriate for real-time applications. It demonstrates the ability to detect arc faults across various scenarios, boosting its practical importance for stakeholders in safety-critical industries.
@@ 1 Introduction
An arcing fault is a particular type of fault that occurs in medium voltage (MV) distribution lines [7]. This fault has the ability to create a shock danger, induce equipment failure, and start a catastrophic fire. Detecting an arcing fault in distribution lines is particularly tough because of its short duration. Thus, to guarantee electrical safety, it’s critical to precisely identify arc faults. However, because of their features, particularly their great unpredictability and strong concealment, it is difficult to discover arc faults in real-world engineering applications.
@@
The occurrence of an arc is directly associated with various types of faults in MV distribution lines[7]. It is usually accepted that an incipient fault includes an arc, particularly an incipient fault in an underground cable [8]. The magnitude of the electric current in these arcing faults could reach several kilo-amperes, making them known as high-current arcing faults. Furthermore, the occurrence of a HIF is consistently associated with an arc, wherein the current of the fault typically ranges from just a few tens of amperes [9]. This type of defect is commonly known as a low current arcing fault.
Researchers worldwide have undertaken substantial investigations on the detection of arc faults that mostly concentrated on identifying arc faults in low-voltage systems [10]. Investigations have been conducted to differentiate between AC and DC configurations in the field of low-voltage systems. However, it is crucial to highlight the insufficient attention given to medium voltage systems in the current literature.
Presently, there exist two basic strategies for the detection of arc faults: traditional approaches [11, 12] and those based on artificial intelligence (AI) [13]. The principal techniques for detecting arc faults involve the analysis of the time domain [11], frequency domain[14], time-frequency domain [15], or the observation of physical phenomena such as sound, light, or electromagnetic radiation. In the past few years, there has been a developing interest in the application of AI in the field of arc fault detection. Numerous AI strategies have been utilized for the goal of identifying arc faults, revealing a varied range of techniques within this field. Within AI methodologies, arc fault detection is classified into two different categories: techniques based on machine learning and neural network-based techniques [13].
In [12], current profiles are examined utilizing Fast Fourier Transform (FFT), and detectors are engineered to focus on the third, fifth, and seventh harmonic components. This approach utilises two criteria to identify and indicate the occurrence of HIAF. An Empirical Mode Decomposition (EMD)-based method is employed for the detection and classification of high-impedance arc faults through the analysis of arc voltage data. In conjunction with EMD, machine learning methods such Artificial Neural Networks (ANN), Support Vector Machines (SVM), and k-Nearest Neighbours (KNN) are utilised. The arc signals undergo processing using EMD to derive Intrinsic Mode Functions (IMFs), which subsequently serve as inputs for the machine learning models utilised in categorisation[16]. An approach for detecting arc fault is proposed that includes SVM, enhanced multi-scale fuzzy entropy (IMFE), and variational mode decomposition (VMD) approaches [17].
Nevertheless, the classic approaches are confronted with the challenges of excessive computational complexity and poor generalization ability when applying the Fast Fourier transform (FFT), Discrete Wavelet Transform (DWT), and similar algorithms for rapid calculations. These approaches offer detection using simple circuits and algorithms, however they are subject to switching noise and load change. In complex power generation and distribution systems, wherein many loads operate at the same time, the chance of arcing in a branch is relatively great. However, it is impractical to place arc fault detection equipment in each branch.Consequently, extracting arc fault features related to the main circuit current presents a major obstacle in the domain of arc fault identification. Studies on the application of machine learning to address the issue of arc fault detection are represented in the target identification of fault features [18] or by building up a classification model to differentiate fault features. Basic classification models include the clustering method and mapping algorithm. These algorithms are less resilient on isolated points and highly disturbed data [19]. The problem of sensitivity to hyperparameter choice bears enormous significance. Machine learning faces a significant problem concerning the potential overfitting of algorithms to the training data [20]. Comprehending these limitations is crucial for stakeholders in safety-critical industries, as it guides the development and application of dependable and effective arc fault detection methods.
@@
The HAVOK technique has been widely implemented across numerous fields, displaying its versatility in electrical engineering [24], and physics[25]. In [24], the HAVOK method is used to analyze real-time streaming data from electrical distribution networks’ distribution-level phasor measurement units. Time series data from behavioral and psychological research[26] can have substantial nonlinear components, making analysis challenging. To address these difficulties, HAVOK analysis is applied as a flexible framework for capturing and analysing underlying dynamics. By offering a linear framework, recreating attractor manifolds, and estimating the amount of nonlinear forcing in the dynamic system, it makes a more complex analysis possible. The significant contributions of this research include:
@@
    This study involves the implementation of a physics-informed HAVOK method for high-impedance arc fault detection, resulting in a detection time that is 50% faster than the methods available in the literature.
@@
    The technique successfully differentiates arc faults from other disturbances in medium voltage distribution lines,achieving an accuracy of 100%, ensuring precise and reliable arc fault identification.
@@
    By providing rapid detection of arc faults within 0.45 ms, the approach enhances the protective mechanisms of electrical distribution networks, potentially preventing major faults and damage.
This article has been arranged as follows. Section 2 discusses the development of the arc model in PSCAD software and generating data for a medium voltage distribution system. Section 3 discusses the approach utilized for arc fault detection.Section 4 presents the experimental scenarios are used to validate the effectiveness of the suggested detection algorithm. Section 5 offers an analysis of our proposed method compared to existing approaches presented in the literature. Section 6 gives the conclusion.
@@ 2 Arc Modelling
A Medium-voltage distribution system has been considered that operates at 12 kV and 50 Hz. The simulation was performed in PSCAD software. The system consists of three buses labeled as B1, B2, and B3, and two loads of 10 MW each shown in Figure 1. An arc model has been integrated into the system to precisely replicate arc faults. A HIAF model is commonly created by connecting a variable resistor, $R_{arc}$, to represent the arc, and a constant resistor to indicate the poor conductivity of the grounding material , $R_{T}$ as shown in Figure 2 [7]. Typically, the nonlinear aspects of the arcing process could be represented by $R_{arc}$, and the current amplitude magnitudes are obtained by modifying $R_{T}$. The $R_{T}$ value impacts the amount of current that can flow through the circuit during an arc fault. A greater $R_{T}$ value often results in reduced current flow, which can influence the detecting capabilities of the system. The sensitivity of the detection system to discover arc faults can be modified by the $R_{T}$ value. In simulations, modifying the $R_{T}$ helps to explore how varying resistance levels affect the system’s response to arc faults. This helps in improving detection method and analysing the dynamics of arc faults.
@@
The nonlinear nature of arc faults could be changed by three fundamental arc characteristics, each providing a distinct physical interpretation and a means for effectively modifying arc resistance. These characteristics play a key role in determining and characterizing residual power. Consequently, they enable the regulation of three components of current distortions, hence boosting simulation accuracy. Ideally, an ideal current distortion should be nearly centered around the zero point and exhibit symmetry. However, in practical settings, distortion offsets occur relative to higher or lower zero points due to variations in the lag between the dissipated power of the arcs and the energizing power. Consequently, the three primary variables in the current waveforms, as illustrated in Figure 3, are the duration, extent, and offset of distortions. Figure 3 includes additional labels to visually depict the three features in a more understandable manner.
@@
$P_{\text{res}}^{(t_{m})}(t)=\begin{cases}a_{1}\left(t-t_{m}-\frac{\text{DURATION}}{2}\right)-b_{1},&t\in\left[t_{m}-\frac{\text{DURATION}}{2},t_{m}+\frac{\text{DURATION}}{2}\right)\\[8.0pt] a_{2}\left(t-t_{m}-\frac{\text{DURATION}}{2}\right)-b_{2},&t\in\left[t_{m}+\frac{\text{DURATION}}{2},t_{m}+\frac{Time}{2}-\frac{\text{DURATION}}{2}\right)\end{cases}$ (8)
@@
OFFSET is to control the distortion offset. The moment at which the maximum arc resistance occurs $t_{m}$ is controlled by OFFSET and satisfies:
@@
where, $R_{\text{arc},0}=R_{\text{arc}}(t_{0})$, and indicates the minimal resistance during $[t_{m}-Time/2,t_{m}]$. $R_{\text{arc},0}$ is the resistance when an arc is highly conductive and which could be ignored in comparison to the huge grounding resistance $R_{T}$Thus, put $R_{\text{arc},0}=1$ and 14 represent as:
@@
$=\left(\frac{Time}{8}-\frac{\text{DURATION}}{4}\right)\cdot m\cdot b_{1}+\frac{1}{2}\cdot\frac{\text{DURATION}}{2}\cdot b_{1}$
$=\left(\frac{Time}{8}-\frac{\text{DURATION}}{4}\right)\cdot m\cdot b_{1}+\ln\text{EXTENT}$
Hence from the equation above, the maximum resistance that arises at $t_{m}$ can be individually reflected by EXTENT if $m$ is taken as 0.
DURATION: Distortion duration is determined by the length of ZOI, which is exactly equal to DURATION. Hence the distorted duration also becomes individually controlled As a result, independent control of all three arc characteristics theoretically make it more flexible in controlling distortion characteristics and achieving the different waveforms.In summary, the modelling of an arc fault entails a systematic process starting with the identification of the time instance $t_{m}$ via the OFFSET parameter, which indicates the initial appearance of arc fault characteristics. Utilising this $t_{m}$, the residual power $P_{\text{res}}(t_{m})$ is computed based on DURATION and EXTENT, capturing the temporal and magnitude aspect of the arc fault. The residual power is utilised to calculate the arc resistance $R_{\text{arc}}$, so establishing a correlation between power dissipation and resistance during the fault. Ultimately, by regulating $R_{\text{arc}}$, arc current waveforms are generated, enabling a realistic simulation of arc faults for analysis and validation. The arc modelling is given in Figure 4.
@@ 3.1 Arc Fault Detection using Hankel Alternative View of Koopman Analysis
The delay embedding approach is utilized to reconstruct the state space of a dynamical system from a time series. The basic concept is to produce a multi-dimensional representation of the time series data, which represents the temporal dependencies and dynamics of the system. In our work of arc fault detection, the time-series current data $i(t)$ taken from system Figure 1, the time shifted measurement of $i(t)$ is utilized and stack them to generate the Hankel matrix $\mathbb{H}_{m}$, as explained in [31]. Subsequently, eigen time delay coordinates are acquired using the method of singular value decomposition (SVD) applied to the Hankel matrix $\mathbb{H}_{m}$.
@@
The columns of equaton 16, and hence equation 17, are efficiently approximated by the first $r$ columns of $U$, making these eigen-time-series a Koopman-invariant measurement system. At the same time, the first $r$ columns of $V$ constitute a time series that captures the magnitude of each column of $U\Sigma$ within the data. A plot of matrix $V$’s first three columns yields an embedded attractor. Then we identify the optimum singular value hard threshold coefficient and computed r. Then the time history of V components are taken and build a regression model on those eigen time delay coordinates.A linear time series that substantially mirrors the input time series is produced by the initial $(r-1)$ variables of $V$ that gives a very nice regression fit. Rather than developing a closed linear model for the initial r variables in V, a linear model is established based on the first $(r-1)$ variables and the last variable $v_{r}$ gives a bad regression fit which will function as a forcing term.
@@
Here $\mathbf{v}=\begin{bmatrix}v_{1}&v_{2}&\cdots&v_{r-1}\end{bmatrix}^{T}$ is a vector of the first $r-1$ eigen-time-delay coordinates. $v_{r}$ serves as an input forcing that influences the linear dynamics in 18, resulting in the nonlinear dynamics in LABEL:eq:_1. Whenever a nonlinearity exists in the system, it undergoes a bursting phenomenon, which serves as an intermittent forcing signature that is forcing operator. The matrix A defines the linear dynamics of the system. It defines the evolution of v(t) without external influences. The matrix B serves as a connection between v(t) and the external force $v_{r}$(t). It describes how the external input (forcing) impacts the evolution of the state.
@@
The forcing operator captures the current fluctuations during arc faults. Upon initiating an arc, it produces unique electrical signatures characterized by abrupt fluctuations in current intensity and high-frequency noise. The presence of an arc fault is indicated by the emergence of these intermittent forcing bursts, detected using the Forcing Operator. The operator enables the detection of these deviations from the norm. Finally, the forcing operator reflecting the nonlinear characteristics of the arc fault in the current data is provided by the $r^{th}$ column. The forcing operator is the last eigentime delay coordinate that is the $r^{th}$ column produced from the SVD of the Hankel matrix, representing the external driving force influencing the system. It captures external influences on the system, such as faults or disturbances, which are different from the natural dynamics of the system. By setting a threshold on the forcing operator, arc faults could be efficiently detected. Figure 5 provides an overview of the framework proposed in this study.
@@
Figure 5: An overview of the suggested arc fault detection technique; HAVOK analysis framework is utilized for arc fault detection in current measurements. Time shifted current measurement is stacked into a Hankel matrix. Its singular value decomposition provides a hierarchy of eigen time series that collectively form a delay-embedded attractor. A best-fit linear regression model is then generated for the delay coordinates $V$; the linear fit is excellent for the first $r-1$ variables, however, the last coordinate $V_{r}$ is not adequately captured by a linear model. Instead, $V_{r}(t)$ functions as a stochastic forcing input to the first $r-1$ delay coordinates, with few yet significant forcing events corresponding to sudden fluctuations in dynamics that indicate the onset of an arc fault, similar to lobe switching in chaotic systems.

### v2

# Enhanced Rapid Detection of High-impedance Arc Faults in Medium Voltage Electrical Distribution Networks

@@ Abstract
High-impedance arc faults in AC power systems have the potential to lead to catastrophic accidents. However, significant challenges exist in identifying these faults because of the much weaker characteristics and variety when grounded with different surfaces. Previous research has concentrated predominantly on arc fault detection in low-voltage systems, leaving a significant gap in medium-voltage applications. In this work, a novel approach has been developed that enables rapid arc fault detection for medium-voltage distribution lines. In contrast to existing black-box feature-based approaches, the Hankel alternative view of the Koopman (HAVOK) analysis developed from nonlinear dynamics has been applied, which not only offers interpretable features but also opens up new application options in the area of arc fault detection. The method achieves a much faster detection speed in 0.45 ms, 99.36% enhanced compared to harmonic randomness and waveform distortion method, thus making it suitable for real-time applications. It demonstrates the ability to detect arc faults across various scenarios, including different grounding surfaces and levels of system noise, boosting its practical importance for stakeholders in safety-critical industries.
@@ 1 Introduction
An arcing fault is a particular type of fault that occurs in medium voltage (MV) distribution lines [7]. This fault has the ability to create a shock danger, induce equipment failure, and start a catastrophic fire. The occurrence of an arc is directly associated with various types of faults in MV distribution lines [7]. It is usually accepted that an incipient fault includes an arc, particularly an incipient fault in an underground cable [8]. The magnitude of the electric current in these arcing faults could reach several kilo-amperes, making them known as high-current arcing faults. Furthermore, the occurrence of a HIF is consistently associated with an arc, wherein the current of the fault typically ranges from just a few tens of amperes [9]. This type of defect is commonly known as a low-current arcing fault.
Detecting arcing faults in distribution lines is particularly challenging due to their transient nature. To ensure electrical safety, accurate identification of these arc faults is essential. However, in practical engineering applications, arc faults are difficult to detect because of their inherent characteristics, especially their high unpredictability and strong concealment within normal system operations.
The detection of high-impedance arc faults in MV distribution systems poses significant challenges not typically observed in low-voltage (LV) environments. MV systems, due to their elevated operating voltages and currents, can conceal fault signatures, as these may closely resemble normal load transients. Moreover, high-impedance faults produce fault currents that are significantly reduced due to elevated fault impedance, causing these currents to frequently mimic normal currents, hence making detection challenging. Reliable identification is made more difficult by the short length and variety of these errors, as well as the electromagnetic interference and noise that are common in MV settings.
Table 1 highlights the key approaches and constraints discovered in the literature for arc fault detection.As shown in Table 1, existing arc fault detection methods face significant drawbacks, including poor generalization, sensitivity to noise and hyperparameters, high computational cost, and limited capacity to capture nonlinear arc dynamics. In response to these constraints, the proposed HAVOK-based framework utilises Koopman operator theory and delay embedding to extract physically meaningful and dynamically informative features. By bridging the gap between black-box AI models and conventional physics-based techniques, this methodology delivers better resilience, interpretability, and computational efficiency for high-impedance arc fault detection.
@@
The HAVOK technique has been widely implemented across numerous fields, displaying its versatility in electrical engineering [13], and physics [14]. In [13], the HAVOK method is used to analyze real-time streaming data from electrical distribution networks’ distribution-level phasor measurement units. Time series data from behavioral and psychological research[15] can have substantial nonlinear components, making analysis challenging. To address these difficulties, HAVOK analysis is applied as a flexible framework for capturing and analysing underlying dynamics. By offering a linear framework, recreating attractor manifolds, and estimating the amount of nonlinear forcing in the dynamic system, it makes a more complex analysis possible. These considerations drive the development of increasingly sensitive and interpretable techniques, like the HAVOK-based analysis provided in this work, by requiring sophisticated detection techniques that can spot small, nonlinear irregularities within complicated electrical signals.
Table 1: Overview of Arc Fault Detection Strategies and Research Directions
Research Focus Description
System Voltage Level The majority body of literature concentrates on arc fault detection in low-voltage systems, with limited consideration given to medium-voltage networks [16].
Detection Strategies Arc fault detection methodologies are primarily categorised into conventional methods and those employing artificial intelligence (AI) [17, 18, 19].
Analytical Domains Detection methods are based on time-domain [17], frequency-domain [20], time-frequency domain [21], or physical signal features such as sound, light, and electromagnetic radiation.
Traditional Signal Processing Methods Traditional techniques, including FFT-based harmonic analysis, are utilised to extract arc characteristics. Nevertheless, these approaches frequently exhibit sensitivity to switching noise and constrained adaptability to fluctuating operating conditions [18].
AI-Based Techniques Machine learning algorithms, including ANN, SVM, and KNN, have been utilised in conjunction with techniques like as EMD and VMD. Despite their success, these approaches are usually limited by overfitting, sensitivity to hyperparameters, and the need for large training datasets [22, 23, 24, 25].
Physics-Informed Approaches Physics-driven approaches, such as time-frequency analysis and sparse representation, have been presented as interpretable alternatives to data-driven models. However, most fail to adequately capture the nonlinear dynamics inherent in arc behavior [21, 26].
Deployment Conventional detection algorithms demonstrate considerable computing complexity and are inappropriate for real-time applications in complicated distribution networks. It is also unfeasible to install arc fault detectors in every branch, complicating the process of feature extraction from main feeder currents [27].
To overcome the key constraints in existing high-impedance arc fault (HIAF) detection strategies, particularly for medium-voltage (MV) systems, this work proposes a physics-informed, data-driven method using HAVOK analysis, established in nonlinear dynamical systems theory. The study utilises Koopman-based HAVOK analysis to attain precise and comprehensible arc fault identification. In contrast to conventional black-box methodologies, our technique offers transparent characteristics through the forcing operator, hence improving explainability. Moreover, it provides rapid detection times, hence enhancing both precision and comprehension of fault dynamics. The primary novel contributions of this work are:
@@
    Application of HAVOK analysis for arc fault detection in MV distribution networks, enabling a linear interpretation of nonlinear system dynamics —a concept previously unexplored in this domain
@@
    Rapid detection capability, achieving arc fault detection within as little as 0.45 milliseconds, which is at least 99.36% enhanced than existing approaches, considerably enhancing detection speed and system responsiveness.
@@
    Robust performance over multiple fault situations, including variable grounding materials (e.g., dry soil, wet cement), system noise levels, and fault current types (low and high current), all verified by simulation using PSCAD.
  4. 4.
    Discrimination between arcing and non-arcing events utilising a Koopman-invariant forcing operator as a fault signature—offering both accuracy and interpretability, a fundamental drawback in past black-box AI-based techniques.
This article has been arranged as follows. Section 2 discusses the development of the arc model in PSCAD software and generating data for a medium voltage distribution system. Section 3 discusses the approach utilized for arc fault detection.Section 4 presents the experimental scenarios used to validate the effectiveness of the suggested detection algorithm. Section 5 offers an analysis of our proposed method compared to existing approaches presented in the literature. Section 6 gives the conclusion.
@@ 2 Arc Modelling
A Medium-voltage distribution system has been considered that operates at 12 kV and 50 Hz. The simulation was performed in PSCAD software. The system consists of three buses labeled as B1, B2, and B3, and two loads of 10 MW each, shown in Figure 1. An arc model has been integrated into the system to precisely replicate arc faults. A HIAF model is commonly created by connecting a variable resistor, $R_{arc}$, to represent the arc, and a constant resistor to indicate the poor conductivity of the grounding material , $R_{T}$ as shown in Figure 2 [7]. Typically, the nonlinear aspects of the arcing process could be represented by $R_{arc}$, and the current amplitude magnitudes are obtained by modifying $R_{T}$. The $R_{T}$ value impacts the amount of current that can flow through the circuit during an arc fault. A greater $R_{T}$ value often results in reduced current flow, which can influence the detection capabilities of the system. The sensitivity of the detection system to discover arc faults can be modified by the $R_{T}$ value. In simulations, modifying the $R_{T}$ helps to explore how varying resistance levels affect the system’s response to arc faults. This helps in improving the detection method and analyzing the dynamics of arc faults.
@@
The nonlinear nature of arc faults could be changed by three fundamental arc characteristics, each providing a distinct physical interpretation and a means for effectively modifying arc resistance. These characteristics play a key role in determining and characterizing residual power. Consequently, they enable the regulation of three components of current distortions, hence boosting simulation accuracy. Ideally, an ideal current distortion should be nearly centered around the zero point and exhibit symmetry. However, in practical settings, distortion offsets occur relative to higher or lower zero points due to variations in the lag between the dissipated power of the arcs and the energizing power. Consequently, the three primary variables in the current waveforms, as illustrated in Figure 3, are the duration, extent, and offset of distortions. Figure 3 includes additional labels to visually depict the three features more understandably.
@@
$P_{\text{res}}^{(t_{m})}(t)=\begin{cases}a_{1}\left(t-t_{m}-\frac{\text{DURATION}}{2}\right)-b_{1},&t\in\left[t_{m}-\frac{\text{DURATION}}{2},\ t_{m}+\frac{\text{DURATION}}{2}\right)\\[10.0pt] a_{2}\left(t-t_{m}-\frac{\text{DURATION}}{2}\right)-b_{2},&t\in\left[t_{m}+\frac{\text{DURATION}}{2},\ t_{m}+\frac{\text{Time}}{2}-\frac{\text{DURATION}}{2}\right)\end{cases}$ (8)
@@
OFFSET is used to control the distortion offset. The moment at which the maximum arc resistance occurs $t_{m}$ is controlled by OFFSET and satisfies:
@@
where, $R_{\text{arc},0}=R_{\text{arc}}(t_{0})$, and indicates the minimal resistance during $[t_{m}-Time/2,t_{m}]$. $R_{\text{arc},0}$ is the resistance when an arc is highly conductive, and which could be ignored in comparison to the huge grounding resistance $R_{T}$Thus, put $R_{\text{arc},0}=1$ and 14 represent as:
@@
$=\left(\frac{\text{Time}}{8}-\frac{\text{DURATION}}{4}\right)\cdot m\cdot b_{1}+\frac{1}{2}\cdot\frac{\text{DURATION}}{2}\cdot b_{1}$
$=\left(\frac{\text{Time}}{8}-\frac{\text{DURATION}}{4}\right)\cdot m\cdot b_{1}+\ln(\text{EXTENT})$ (12)
Hence, from the equation above, the maximum resistance that arises at $t_{m}$ can be individually reflected by EXTENT if $m$ is taken as 0.
DURATION: Distortion duration is determined by the length of ZOI, which is exactly equal to DURATION. Hence, the distorted duration also becomes individually controlled As a result, independent control of all three arc characteristics theoretically make it more flexible in controlling distortion characteristics and achieving the different waveforms.In summary, the modelling of an arc fault entails a systematic process starting with the identification of the time instance $t_{m}$ via the OFFSET parameter, which indicates the initial appearance of arc fault characteristics. Utilising this $t_{m}$, the residual power $P_{\text{res}}(t_{m})$ is computed based on DURATION and EXTENT, capturing the temporal and magnitude aspect of the arc fault. The residual power is utilised to calculate the arc resistance $R_{\text{arc}}$, so establishing a correlation between power dissipation and resistance during the fault. Ultimately, by regulating $R_{\text{arc}}$, arc current waveforms are generated, enabling a realistic simulation of arc faults for analysis and validation. The arc modelling is given in Figure 4.
@@ 3.1 Arc Fault Detection using Hankel Alternative View of Koopman Analysis
The delay embedding approach is utilized to reconstruct the state space of a dynamical system from a time series. The basic concept is to produce a multi-dimensional representation of the time series data, which represents the temporal dependencies and dynamics of the system. In our work of arc fault detection, the time-series current data $i(t)$ taken from system Figure 1, the time shifted measurement of $i(t)$ is utilized and stack them to generate the Hankel matrix $\mathbb{H}_{m}$, as explained in [32].In our approach, the Hankel matrix is formed by aggregating 40 time-shifted rows of the input current signal, so establishing the embedding dimension (q) at 40. This value provides an effective compromise between adequately recording temporal dynamics and ensuring computational efficiency, as informed by previous HAVOK investigations [33, 32]. An embedding dimension of 40 was selected based on earlier HAVOK studies to represent the system’s dynamic behavior well. In practice, this number should be tuned to the sample rate. A greater sampling rate provides for a larger embedding dimension to resolve finer dynamics, whereas a lower rate may require fewer delays. To guarantee significant reconstruction of the system’s attractor, the total embedding window (embedding dimension × sampling interval) should, as a general rule, span at least one cycle of the main system frequency (20 ms for a 50 Hz system, for example). The embedding dimension q=40 was selected based on a trade-off between capturing critical system dynamics and avoiding excessive model complexity. Similarly, the window length was empirically selected to achieve robust detection performance without representing misleading or non-generalizable fluctuations in the signal[33]. The window length (p) is dynamically established as the residual signal length following delay embedding, optimizing data usage while maintaining matrix stability. These parameters facilitate a robust extraction of Koopman-invariant coordinates while preventing overfitting and minimizing excessive noise. Subsequently, eigen time delay coordinates are acquired using the method of singular value decomposition (SVD) applied to the Hankel matrix $\mathbb{H}_{m}$.
@@
The columns of equation 16, and hence equation 17, are efficiently approximated by the first $r$ columns of $U$, making these eigen-time series a Koopman-invariant measurement system. At the same time, the first $r$ columns of $V$ constitute a time series that captures the magnitude of each column of $U\Sigma$ within the data. A plot of matrix $V$’s first three columns yields an embedded attractor. Then we identify the optimum singular value hard threshold coefficient and compute r. Then the time history of V components is taken, and a regression model on those eigen time delay coordinates. A linear time series that substantially mirrors the input time series is produced by the initial $(r-1)$ variables of $V$ that gives a very nice regression fit. Rather than developing a closed linear model for the initial r variables in V, a linear model is established based on the first $(r-1)$ variables, and the last variable $v_{r}$ gives a bad regression fit, which will function as a forcing term.
@@
Here $\mathbf{v}=\begin{bmatrix}v_{1}&v_{2}&\cdots&v_{r-1}\end{bmatrix}^{T}$ is a vector of the first $r-1$ eigen-time-delay coordinates. $v_{r}$ serves as an input forcing that influences the linear dynamics in 18, resulting in the nonlinear dynamics in 13. Whenever a nonlinearity exists in the system, it undergoes a bursting phenomenon, which serves as an intermittent forcing signature that is a forcing operator. The matrix A defines the linear dynamics of the system. It defines the evolution of v(t) without external influences. The matrix B serves as a connection between v(t) and the external force $v_{r}$(t). It describes how the external input (forcing) impacts the evolution of the state.
@@
The forcing operator captures the current fluctuations during arc faults. Upon initiating an arc, it produces unique electrical signatures characterized by abrupt fluctuations in current intensity and high-frequency noise. The presence of an arc fault is indicated by the emergence of these intermittent forcing bursts, detected using the Forcing Operator. The operator enables the detection of these deviations from the norm. Finally, the forcing operator reflecting the nonlinear characteristics of the arc fault in the current data is provided by the $r^{th}$ column. The forcing operator is the last eigentime delay coordinate that is the $r^{th}$ column produced from the SVD of the Hankel matrix, representing the external driving force influencing the system. It captures external influences on the system, such as faults or disturbances, which are different from the natural dynamics of the system. By setting a threshold on the forcing operator, arc faults could be efficiently detected. Figure 5 provides an overview of the framework proposed in this study. The entire detection procedure is encapsulated in Algorithm 1, which emphasises the formulation of the Hankel matrix, the extraction of forcing coordinates, and the classification of arc faults predicated on the magnitude of the forcing operator.
@@
Figure 5: An overview of the suggested arc fault detection technique; HAVOK analysis framework is utilized for arc fault detection in current measurements. Time-shifted current measurement is stacked into a Hankel matrix. Its singular value decomposition provides a hierarchy of eigen time series that collectively form a delay-embedded attractor. A best-fit linear regression model is then generated for the delay coordinates $V$; the linear fit is excellent for the first $r-1$ variables, however, the last coordinate $V_{r}$ is not adequately captured by a linear model. Instead, $V_{r}(t)$ functions as a stochastic forcing input to the first $r-1$ delay coordinates, with few yet significant forcing events corresponding to sudden fluctuations in dynamics that indicate the onset of an arc fault, similar to lobe switching in chaotic systems.
Algorithm 1 HAVOK-based Arc Fault Detection with Forcing Operator
0:  Time-series data $i(t)$, sampling interval $\Delta t$, number of delay embeddings $stack\_max$
0:  Detection of arc faults via forcing operator threshold
1:  Construct the Hankel matrix $H$ by stacking $stack\_max$ delayed versions of $i(t)$
2:  Perform Singular Value Decomposition: $H=U\Sigma V^{T}$
3:  Determine model rank $r$ using an optimal hard threshold on singular values
4:  Compute derivatives of $V$ using fourth-order central difference method
5:  Define delay-embedded state matrix $x=V[:,1:r]$
6:  Identify the last eigen time-delay coordinate $v_{r}=x[:,r-1]$ as the forcing operator
7:  Detect arc fault onset by identifying bursts in $v_{r}$ beyond predefined thresholds:
8:  if $0.06\leq|v_{r}|\leq 0.18$ then
9:     return Arc fault detected
10:  else if $|v_{r}|>0.2$ then
11:     return Other fault type (non-arc)
12:  else if $|v_{r}|<0.045$ then
13:     return Non-arcing disturbance (e.g., load switching)
14:  else
15:     return Inconclusive or borderline case
16:  end if